### v1

# Optimization of Superconducting Niobium Nitride Thin Films via High-Power Impulse Magnetron Sputtering

Abstract
We report a systematic comparison of niobium nitride thin films deposited on oxidized silicon substrates by reactive DC magnetron sputtering and reactive high-power impulse magnetron sputtering (HiPIMS). After determining the nitrogen gas concentration that produces the highest superconducting critical temperature for each process, we characterize the dependence of the critical temperature on film thickness. The optimal nitrogen concentration is higher for HiPIMS than for DC sputtering, and HiPIMS produces higher critical temperatures for all thicknesses studied. We attribute this to the HiPIMS process enabling the films to get closer to optimal stoichiometry before beginning to form a hexagonal crystal phase that reduces the critical temperature, along with the extra kinetic energy in the HiPIMS process enabling greater adatom mobility and improving crystallinity. We also study the effects of an aluminum nitride buffer layer and substrate heating on the critical temperature.
preprint: AIP/123-QED
I Introduction
Niobium nitride (NbN) is a refractory metal with a superconducting critical temperature ($T_{\mathrm{C}}$) as high as $\approx 16$ K.Shy, Toth, and Somasundaram (1973) NbN thin films are utilized in a wide range of superconducting device applications, including photodetectors such as hot electron bolometersGousev et al. (1994) and superconducting nanowire single-photon detectors;Holzman and Ivry (2019) and Josephson junction devices such as SIS heterodyne mixers,McGrath et al. (1991) RSFQ digital logic,Shao et al. (2024) and superconducting qubits.Kim et al. (2021) Due to their high kinetic inductivity, NbN thin films are of interest in applications such as slow-wave transmission lines,Santavicca et al. (2021) tunable couplers,Colangelo et al. (2021) and travelling wave parametric amplifiers.Adamyan et al. (2016) For many of these applications, very thin films ($\lesssim 10$ nm) are required, and hence a number of groups have explored different techniques for producing high-quality ultra-thin NbN films.Ilin et al. (2008); Espiau de Lamaëstre, Odier, and Villégier (2007); Iovan et al. (2023); Chockalingam et al. (2008); Wright, Xing, and Jena (2023); Schneider et al. (2009); Shiino et al. (2010); Rhazi et al. (2021)
Several crystal phases of NbN can occur under normal conditions. The cubic NaCl-like $\delta$-phase has the highest $T_{\mathrm{C}}$ and occurs when a near-stoichiometric reaction between nitrogen and niobium takes place.Shy, Toth, and Somasundaram (1973) Increasing the nitrogen to a $1{:}1$ stoichiometry can result in the formation of the hexagonal $\epsilon$-phase with a lower $T_{\mathrm{C}}$. On the other hand, reducing the nitrogen concentration can lead to the formation of tetragonal Nb4N3 or hexagonal Nb2N, which also have a lower $T_{\mathrm{C}}$.Benkahoul et al. (2004); Wright, Xing, and Jena (2023) Optimizing the nitrogen concentration is thus important for maximizing the $T_{\mathrm{C}}$ of NbN thin films.
Other factors also affect the film structure and $T_{\mathrm{C}}$, including the choice of substrate and its temperature during deposition. While Si substrates are often desired due to their availability in larger wafer sizes and their compatibility with standard fabrication processes, MgO and sapphire substrates have been shown to produce NbN films with higher $T_{\mathrm{C}}$ due to their improved lattice match with NbN.Miki et al. (2007); Kang et al. (2003); Espiau de Lamaëstre, Odier, and Villégier (2007) The use of an AlN buffer layer has also been shown to improve $T_{\mathrm{C}}$ on substrates such as Si.Shiino et al. (2010); Rhazi et al. (2021)
Heating the substrate can improve the crystallinity of the film, and single-domain epitaxial films can be produced using an appropriate substrate at sufficiently high temperature.Kim et al. (2021); Wright, Xing, and Jena (2023); Iovan et al. (2023); Chockalingam et al. (2008) These high temperatures, however, are not compatible with some materials and fabrication processes, including lift-off processes and standard CMOS processes. Single-domain epitaxial films also have a lower kinetic inductivity than polycrystaline films and hence may not be desirable for applications that utilize the high kinetic inductivity of NbN.
Sputter deposition is the most widely used approach for producing NbN thin films. High-power impulse magnetron sputtering (HiPIMS) is a sputter deposition technique in which a DC power supply is operated in a pulsed mode, resulting in high-intensity plasma excitation during the ${\sim 100}$ $\mu$s pulses.Lundin and Sarakinos (2012); Gaines (2019) This translates to a higher kinetic energy of the sputtered target material relative to conventional DC sputtering. Recent work found an increased $T_{\mathrm{C}}$ in NbN films deposited via HiPIMS as compared to DC sputtering.Kalal et al. (2022) This work studied relatively thick films deposited on a Si substrate with a TiN buffer layer. Here, we report a systematic comparison of DC sputtered and HiPIMS films deposited on oxidized silicon substrates, including the dependence of $T_{\mathrm{C}}$ on the nitrogen gas concentration during deposition and the film thickness. We also study the use of an AlN buffer layer and substrate heating with the HiPIMS process.
II Deposition Parameters
All depositions were performed in a Kurt J. Lesker PRO Line PVD 75 system with $3$-inch diameter sputter targets. The system is pumped by a turbomolecular pump backed by an oil-free scroll pump and the chamber has a base pressure of $<1\times 10^{-7}$ torr. The system uses a sputter-up configuration with a vertical distance of approximately $12$ cm between the targets and the sample stage. The sample stage was rotated at $10$ rpm during all depositions. The sample stage can be heated to $350$ C using a quartz lamp heater.
NbN films were deposited using a Nb target in a reactive process with nitrogen gas. The Nb target was specified as 99.95% pure (excluding tantalum). Ultra high purity argon gas (99.9999% purity) and ultra high purity nitrogen gas (99.999% purity) were introduced into the chamber via separate mass flow controllers. The chamber was pumped to $\lesssim 5\times 10^{-7}$ torr prior to deposition. During all depositions, a gate valve is used to throttle the turbomolecular pump and the mass flow controllers set the flow rates to achieve a pressure of $3.0\times 10^{-3}$ torr. The nitrogen gas concentration is specified as a percentage of the argon gas flow rate (not as a percentage of the total gas flow). For example, an argon flow rate of $12.5$ sccm and a nitrogen flow rate of $2.5$ sccm would be specified as a $20$% nitrogen gas concentration.
Depositions were performed using both DC sputtering and high-power impulse magnetron sputtering (HiPIMS). The HiPIMS depositions utilized an Impulse $2$ kW pulsed power module from Starfire Industries. This module allows for a user-configurable voltage pulse width, height, and repetition rate. All HiPIMS depositions described here used a $1$ kHz repetition rate and a $100$ $\mu$s pulse width. The HiPIMS was performed in a power-limited mode set to a time-average power of $250$ W; this determined the resulting voltage pulse height, which had a value of approximately $480$ V. The main pulse was followed by a positive kick pulse with an amplitude of approximately $+50$ V and a duration of $50$ $\mu$s. This kick pulse has been shown to increase the deposition rate by driving ionized target material to the sample, as the high peak power of the HiPIMS process can result in a higher ionization fraction of the target material.Gaines (2019); Kalal et al. (2022) An example of the resulting voltage and current pulses is shown in figure 1. The peak power, which occurs at the maximum of the current pulse, is approximately $3$ kW.
Figure 1: Voltage and current pulses used in the HiPIMS process. The pulses correspond to a time-average power of $250$ W and a peak power of approximately $3$ kW.
The NbN films deposited with DC sputtering used a power of $150$ W. The different time-average powers for the DC and HiPIMS depositions were chosen to achieve a similar deposition rate of $\approx 1$/s, although the deposition rate decreased as the concentration of nitrogen gas was increased, as discussed in the next section.
[…]
III Comparison of DC and HiPIMS Films
NbN films were deposited onto unheated substrates at varying nitrogen gas concentrations using both DC sputtering and HiPIMS. The deposition rate varies with the nitrogen gas concentration, so the deposition rate had to be determined for each nitrogen concentration. This was done by performing a $900$ s deposition, scribing and snapping the resulting sample, and imaging the sample’s cross-section in a scanning electron microscope (SEM) to determine the film thickness. An example SEM image of a $120$ nm thick film deposited via HiPIMS at $10\%$ nitrogen concentration is shown in figure 2.
Figure 2: Scanning electron micrograph of cross-section of 120 nm thick NbN film deposited via HiPIMS at $10\%$ nitrogen gas concentration.
The resulting deposition rates are shown in figure 3. After these were determined, $120$ nm thick films were deposited at different nitrogen gas concentrations using both DC sputtering and HiPIMS, and the $T_{\mathrm{C}}$ of each film was measured. Here we define $T_{\mathrm{C}}$ as the temperature at which the resistance drops to half of its normal-state value at 18 K. Figure 3 shows these results. The nitrogen concentration yielding the highest $T_{\mathrm{C}}$ was 15% for DC sputtering and 20% for HiPIMS.
Figure 3: (a) $T_{\mathrm{C}}$ of $120$ nm thick NbN films deposited via HiPIMS and DC sputtering onto unheated oxidized silicon wafers with different nitrogen gas concentrations. (b) Deposition rate as a function of nitrogen gas concentration. DC sputtering was performed at $150$ W and HiPIMS was performed at a time-average power of $250$ W.
X-ray diffraction (XRD) was used to determine the lattice constant of films deposited via DC sputtering and HiPIMS at the optimal nitrogen concentration for each process. To increase the XRD signal, thicker films of $1$ $\mu$m were grown for this characterization. The results are shown in figure 4. Peaks consistent with a cubic crystal structure and corresponding to Miller indices ($hkl$) = (111) and (200) are observed. For a simple cubic structure, the lattice constant $a$ is related to the interplane spacing $d$ by $a=d\sqrt{h^{2}+k^{2}+l^{2}}$. The diffraction angle $\theta$ is related to $d$ via Bragg’s law, $n\lambda=2d\sin\theta$, where here we have first-order diffraction ($n=1$). The XRD system uses a Cu Kα x-ray source, which has a wavelength $\lambda=1.5406$.
For each XRD peak in figure 4, a Gaussian fit was performed to extract the peak center. This angle was then used to calculate the associated lattice constant $a$. The results are given in table 1. We find a lattice constant of $4.40$ for the HiPIMS film and a smaller lattice constant of $4.38$ for the DC sputtered film. Previous work has found that an increase in nitrogen concentration is associated with an increase in the lattice constant.Shy, Toth, and Somasundaram (1973); Kalal, Gupta, and Rawat (2021); Torche et al. (2003) The larger lattice constant in the HiIPMS film is consistent with the picture of the HiPIMS process enabling us to get closer to optimal stoichiometry before beginning to form the $\epsilon$-NbN crystal phase that reduces $T_{\mathrm{C}}$.
Figure 4: X-ray diffraction of $1$ $\mu$m thick films deposited via DC sputtering and HiPIMS using the optimal nitrogen concentration for each process, which was 15% for DC sputtering and 20% for HiIPMS.
Table 1: Cubic lattice constant values calculated from XRD data in figure 4 assuming a simple cubic crystal structure and assigning Miller indices of (111) to the peak at $35^{\circ}$ and Miller indices of (200) to the peak at $41^{\circ}$.
HiPIMS film DC sputtered film
$2\theta$ peak lattice constant $2\theta$ peak lattice constant
$35.323$ $4.3976$ $35.493$ $4.3772$
$41.027$ $4.3964$ $41.141$ $4.3847$
Figure 5: $T_{\mathrm{C}}$ of NbN films deposited via HiPIMS and DC sputtering as a function of film thickness. Films were deposited onto unheated silicon substrates with a native oxide using the optimal nitrogen concentration for each deposition technique.
Using the optimal nitrogen concentration for the DC and HiPIMS processes respectively, a set of films of varying thickness between $6$ nm and $120$ nm were deposited to explore the dependence of $T_{\mathrm{C}}$ on the film thickness. The results are shown in figure 5. We see that the HiPIMS films have a higher $T_{\mathrm{C}}$ than the DC sputtered films at all thicknesses studied, with the difference becoming larger for the thinnest samples.
IV Optimization of HiPIMS Films
To study the potential for improvement of the $T_{\mathrm{C}}$ of our HiPIMS films, we deposited a set of films of different thicknesses on top of a $20$ nm AlN buffer layer, and on top of a heated substrate with a $40$ nm AlN buffer layer. The AlN layer was deposited as described previously, and the the NbN was then deposited in the same deposition system without breaking vacuum. All NbN films were deposited using HiPIMS with a $20$% nitrogen gas concentration. The resulting critical temperatures are shown in figure 6. As has been seen in previous work with DC sputtering, the use of the AlN buffer layer increased $T_{\mathrm{C}}$. This is attributed to lattice matching between the AlN and cubic NbN.Shiino et al. (2010); Rhazi et al. (2021)
We also investigated the effect of depositing onto heated substrates. Initially, the quartz lamp heater in the deposition system was powered with an AC power supply. This produced an instability in the HiPIMS process, with the plasma visibly flickering at the power line frequency. Powering the heater with a DC power supply eliminated this instability. All heated depositions presented here used the DC power supply.
First, we deposited NbN onto bare substrates (Si with a native oxide and no AlN buffer layer) that were heated to $350$ C. The resistance versus temperature curves of the resulting films showed a reduced and significantly broadened superconducting transition ($\Delta T_{\mathrm{C}}>1$ K). We believe that this resulted from an unwanted chemical interaction between the substrate and the NbN enabled by the combination of thermal energy and the high kinetic energy produced by the HiPIMS pulses. We found that this could be prevented and sharp transitions could be obtained by first depositing the AlN layer at room temperature, then heating the substrate to $350$ C, and then depositing the NbN. For these samples, the thickness of the AlN buffer layer was increased to $40$ nm to ensure that it was effective at inhibiting the interaction between the NbN and the underlying substrate. With the buffer layer, the heated NbN films showed a further increase in $T_{\mathrm{C}}$, as seen in figure 6. This is consistent with the expectation that a higher substrate temperature results in greater adatom mobility during the deposition process, which improves crystallization as local nitrogen excesses and deficits in the lattice can diffuse out and long-range crystalline order can develop.Kalal, Gupta, and Rawat (2021); Thakoor et al. (1985)
[…]
In each case, the HiPIMS process produced films with a lower sheet resistance and a higher $T_{\mathrm{C}}$ than DC sputtering. Interestingly, and in contrast to previous results,Shiino et al. (2010) we observe an increase in the sheet resistance when the film is deposited on an AlN buffer layer as compared to the bare substrate. We speculate that this may arise from the surface roughness of the sputtered AlN. If this is the case, one would expect to see this effect reduced in thicker NbN films. To test this, we deposited two $50$ nm thick NbN films via HiPIMS at $20$% nitrogen gas concentration, one with a $40$ nm AlN buffer layer and one without the buffer layer. These films both had the same sheet resistance, $37$ $\Omega$/sq, at room temperature. This is consistent with the idea that the higher sheet resistance of the $10$ nm NbN films on AlN is related to AlN surface roughness. In future work, it may be useful to explore approaches for improving the surface roughness of the sputtered AlN buffer layer.
Figure 6: $T_{\mathrm{C}}$ as a function of thickness of NbN films deposited via HiPIMS with 20% nitrogen concentration. Deposition onto unheated silicon substrates with a native oxide is compared to films deposited on the same substrate with the addition of a $20$ nm thick AlN buffer layer, as well as films with a $40$ nm thick AlN buffer layer that were heated to $350$ C for the NbN deposition.
Table 2: Comparison of 10 nm thick NbN films
Deposition type Sheet resistance at 296 K Sheet resistance at 18 K Critical temperature
DC $652$ $\Omega$/sq $908$ $\Omega$/sq $5.90$ K
HiPIMS $281$ $\Omega$/sq $328$ $\Omega$/sq $9.30$ K
DC with AlN buffer 921 $\Omega$/sq 1,613 $\Omega$/sq 7.25 K
HiPIMS with AlN buffer $346$ $\Omega$/sq $419$ $\Omega$/sq $11.41$ K
heated DC with AlN buffer 691 $\Omega$/sq 1,129 $\Omega$/sq 8.90 K
heated HiPIMS with AlN buffer $323$ $\Omega$/sq 391 $\Omega$/sq $12.55$ K
V Conclusions
In reactive DC sputtered NbN films, it has been shown that a hexagonal crystal phase with a lower $T_{\mathrm{C}}$ begins to form as one approaches a $1{:}1$ stoichiometric ratio.Benkahoul et al. (2004); Wright, Xing, and Jena (2023) We have found that, for reactive HiPIMS, the nitrogen gas concentration that produces the highest $T_{\mathrm{C}}$ is greater than that for DC sputtering, and the maximum $T_{\mathrm{C}}$ for the HiPIMS films is greater than the maximum $T_{\mathrm{C}}$ for the DC sputtered films. This holds true across all film thicknesses studied, from $6$ nm to $120$ nm. We believe that this is the result of the HiPIMS process enabling the films to get closer to optimal stoichiometry before beginning to form the hexagonal $\epsilon$-NbN crystal phase. The lower normal-state resistivity of the HiPIMS films is attributed to increased adatom mobility, and hence improved film crystallinity, enabled by the higher kinetic energy in the HiPIMS process.
We found that the $T_{\mathrm{C}}$ of HiPIMS films can be increased through the use of an AlN buffer layer, and by depositing an AlN buffer layer and then subsequently heating the substrate to $350$ C. In all deposition scenarios tested, the HiPIMS films have a higher $T_{\mathrm{C}}$ and a lower normal-state resistivity than the DC sputtered films.
The HiPIMS deposition process presents a large parameter space, and further exploration of this space may be productive. For example, we have not yet studied the effect of varying parameters such as the deposition chamber pressure and the height and duration of the HiPIMS voltage pulses. In future work, we also plan to explore the benefits of the HiPIMS process for other materials in the superconducting transition metal nitride family.
Acknowledgements.
This work was supported by the National Science Foundation through awards ECCS-2117007 and ECCS-2000778.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.